# Replacing Paths with Connection-Biased Attention for Knowledge Graph Completion

### v1

@@ Discussion
However, we argue that omitting the use of paths has its strengths. Firstly, computing paths between target pairs on the fly is highly expensive and researchers often fix a path length and precompute a set of paths between target pairs of test set. But, this can be a limitation in real applications, as one will come across a new target entity pair during test time and has to compute paths between them. This also comes with an extra set of hyperparameter tuning regarding how many paths to use, what should be the path length, how to represent inverse relations, and how to aggregate each path.
In contrast to these, our proposed method inflicts only a handful of new learnable bias vectors for each connection type between tokens. The entity-role introduced in our paper is also an intuitive construct and adds negligible overhead to represent all entities in the model. Since these two constructs capture the similarity of involved entities, the model implicitly learns relative distance information along with possible path information between tokens.
@@
We propose CBLiP, a KG link prediction model using inexpensive and intuitive use of entity-role and connection-bias in a subgraph. We show the effectiveness of our model in two KG settings in different tasks where CBLiP showcases excellent performance while being intuitive and relatively simpler to its contemporaries.
We identify two future directions from this work. KGs are growing with the inclusion of newer facts and sometimes the new facts added could have relations that are unseen during training as well. This calls for models focusing on graph topology to reason effectively in the presence of unseen entities and relations. We also notice that the graph-connectivity property in the standard benchmark datasets varies to a large degree. A study to investigate the impact of such properties would greatly benefit the field of knowledge graph based reasoning.
Appendix A Supplementary Materials
In this supplementary material, we provide the URL to our model CBLiP’s codebase, the dataset statistics in inductive and transductive settings, and additional transductive relation prediction results.
Appendix B Codebase
We provide the pytorch implementation of our code and hyperparameter configurations for all our experiments at: https://anonymous.4open.science/r/CBLiP-5293/.
Appendix C Inductive Entity Prediction
Dataset
@@ Appendix C Inductive Entity Prediction
Version Split WN18RR FB15k-237 NELL-995
$||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2 $||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2 $||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2
v1 train 9 2746 5410 630 183 2000 4245 489 14 10915 4687 414
test 9 922 1618 188 146 1500 1993 205 14 225 833 100
v2 train 10 6954 15262 1838 203 3000 9739 1166 88 2564 8219 922
test 10 2923 4011 441 176 2000 4145 478 79 4937 4586 476
v3 train 11 12078 25901 3097 218 4000 17986 2194 142 4647 16393 1851
test 11 5084 6327 605 187 3000 7406 865 122 4921 8048 809
v4 train 9 3861 7940 934 222 5000 27203 3352 77 4922 7546 876
test 9 7208 12334 1429 204 3500 11714 1424 61 3294 7073 731
Appendix D Transductive Relation Prediction
Datasets
Initial KG experiments were done on Freebase and Wordnet datasets, FB15k and WN18, whose test sets contained inverse triples of training triples. It caused simpler models to perform well in test by simply memorizing training data. (Toutanova and Chen 2015) proposed a corrected versions of these datasets- FB15K-237 and WN18RR, respectively. We have presented results on FB15k-237, WN18RR, and NELL995 in the main text of our paper.
@@ Appendix D Transductive Relation Prediction
Table 6: Statistics of transductive KG datasets
WN18 WN18RR FB15K FB15K-237 NELL995 DDB14
#nodes 40,943 40,943 14,951 14,541 63,917 9,203
#relations 18 11 1345 237 198 14
#training 141,442 86,835 483,142 272,115 137,465 36,561
#validation 5,000 2,824 50,000 17,526 3,907 3,897
#test 5,000 2,924 59,071 20,438 3,964 3,882
Experimental Evaluation
We have presented the experimental evaluation on widely used FB15k-237, WN18RR and NELL995 in our main text. Here, we present results in the other 3 datasets. Table 7 shows CBLiP’s performance in comparison with the relation prediction baseline models. CBLiP performs achieves competitive performance in WN18 and DDB14. It struggles in FB15K. We think the high triples to entities ratio in this dataset could be the reason for this. The dense neighborhood information is best captured with a path-based model like PathCon (Wang, Ren, and Leskovec 2021).
Table 7: Transductive relation prediction in KG datasets; Bold and underlined text represent best results and 2nd best results, respectively.
FB15K WN18 DDB14
Methods MRR Hits@1 Hits@3 MRR Hits@1 Hits@3 MRR Hits@1 Hits@3
TransE 0.962 0.940 0.982 0.971 0.955 0.984 0.966 0.948 0.980
ComplEx 0.901 0.844 0.952 0.985 0.979 0.991 0.953 0.931 0.968
DistMult 0.661 0.439 0.868 0.786 0.584 0.987 0.927 0.886 0.961
RotatE 0.979 0.967 0.986 0.984 0.979 0.986 0.953 0.934 0.964
SimplE 0.983 0.972 0.991 0.972 0.964 0.976 0.924 0.892 0.948
QuatE 0.983 0.972 0.991 0.981 0.975 0.983 0.946 0.922 0.962
DRUM 0.945 0.945 0.978 0.969 0.956 0.98 0.958 0.930 0.987
PathCon 0.984 0.974 0.995 0.993 0.988 0.998 0.980 0.966 0.995
CBLiP 0.863 0.763 0.962 0.991 0.985 0.996 0.981 0.967 0.995

### v3

@@ Discussion
However, we argue that omitting the use of paths has its strengths. Firstly, computing paths between target pairs on the fly is highly expensive and researchers often fix a path length and precompute a set of paths between target pairs of test set. However, this can be a limitation in real applications, as one will come across a new target entity pair during test time and has to compute paths between them. This also comes with an extra set of hyperparameter tuning regarding path length, how many paths to use, how to represent inverse relations, and how to aggregate each path.
In contrast to these, our proposed method adds only a handful of new learnable bias vectors for each connection type between tokens. Adding the connection bias to the Transformer encoder block does not add significant computational overhead since these bias computations do not require additional matrix multiplications. The scaled dot product attention module (that we modify by adding connection-bias vectors) retains its complexity of $\mathcal{O}(N^{2}d)$ for a sequence of $N$ tokens with $d$ feature dimension.
The entity role introduced in our paper is also an intuitive construct and adds negligible overhead to represent all entities in the model. Since these two constructs capture the similarity of involved entities, the model implicitly learns relative distance information along with possible path information between tokens.
@@
We propose CBLiP, a KG link prediction model using inexpensive and intuitive use of entity role and connection-bias in a subgraph. We show the effectiveness of our model in two KG settings in different tasks where CBLiP showcases excellent performance while being intuitive and relatively simpler to its contemporaries. Future work can address the fully inductive setting, where entities and relations may be seen at test time that were not present at training time.
A. Codebase
We provide the pytorch implementation of our code and hyperparameter configurations for all our experiments at: https://github.com/shoron-dutta/CBLiP.
B. Inductive Entity Prediction
B1. Hyperparameter Selection
B2. Dataset
@@ B. Inductive Entity Prediction
Table 6: WN18RR
Version Split $||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2
v1 train 9 2746 5410 630
test 9 922 1618 188
v2 train 10 6954 15262 1838
test 10 2923 4011 441
v3 train 11 12078 25901 3097
test 11 5084 6327 605
v4 train 9 3861 7940 934
test 9 7208 12334 1429
Table 7: FB15k-237
Version Split $||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2
v1 train 183 2000 4245 489
test 146 1500 1993 205
v2 train 203 3000 9739 1166
test 176 2000 4145 478
v3 train 218 4000 17986 2194
test 187 3000 7406 865
v4 train 222 5000 27203 3352
test 204 3500 11714 1424
Table 8: NELL-995
Version Split $||\mathcal{R}|$$|\mathcal{E}|$TR1 #TR2
v1 train 14 10915 4687 414
test 14 225 833 100
v2 train 88 2564 8219 922
test 79 4937 4586 476
v3 train 142 4647 16393 1851
test 122 4921 8048 809
v4 train 77 4922 7546 876
test 61 3294 7073 731
Table 9: Statistics of transductive
KG datasets
WN18 WN18RR
#nodes 40,943 40,943
#relations 18 11
#training 141,442 86,835
#validation 5,000 2,824
#test 5,000 2,924
FB15K FB15K-237
#nodes 14,951 14,541
#relations 1,345 237
#training 483,142 272,115
#validation 50,000 17,526
#test 59,071 20,438
NELL995 DDB14
#nodes 63,917 9,203
#relations 198 14
#training 137,465 36,561
#validation 3,907 3,897
#test 3,964 3,882
@@
Table 10: Transductive relation prediction in KG datasets
Methods MRR Hits@1 Hits@3
FB15K
TransE 0.962 0.940 0.982
ComplEx 0.901 0.844 0.952
DistMult 0.661 0.439 0.868
RotatE 0.979 0.967 0.986
SimplE 0.983 0.972 0.991
QuatE 0.983 0.972 0.991
DRUM 0.945 0.945 0.978
PathCon 0.984 0.974 0.995
CBLiP 0.863 0.763 0.962
WN18
TransE 0.971 0.955 0.984
ComplEx 0.985 0.979 0.991
DistMult 0.786 0.584 0.987
RotatE 0.984 0.979 0.986
SimplE 0.972 0.964 0.976
QuatE 0.981 0.975 0.983
DRUM 0.969 0.956 0.980
PathCon 0.993 0.988 0.998
CBLiP 0.991 0.985 0.996
DDB14
TransE 0.966 0.948 0.980
ComplEx 0.953 0.931 0.968
DistMult 0.927 0.886 0.961
RotatE 0.953 0.934 0.964
SimplE 0.924 0.892 0.948
QuatE 0.946 0.922 0.962
DRUM 0.958 0.930 0.987
PathCon 0.980 0.966 0.995
CBLiP 0.981 0.967 0.995
C. Transductive Relation Prediction
C1. Datasets
Initial KG experiments were done on Freebase and Wordnet datasets, FB15k and WN18, whose test sets contained inverse triples of training triples. It caused simpler models to perform well during test by memorizing training data. (Toutanova and Chen 2015) proposed a corrected versions of these datasets- FB15K-237 and WN18RR, respectively. We have presented results on FB15k-237, WN18RR, and NELL995 in the main text of our paper.
@@ C. Transductive Relation Prediction
C2. Experimental Evaluation
Here, we present results in the other 3 datasets. Table LABEL:tab:transductive_results_part_2 shows CBLiP’s performance in comparison with the relation prediction baseline models. CBLiP performs achieves competitive performance in WN18 and DDB14. It struggles in FB15K. We think the high triples to entities ratio in this dataset could be the reason for this. The dense neighborhood information is best captured with a path-based model like PathCon (Wang, Ren, and Leskovec 2021).